Characteristise the flow of liquid which is flowing through a tube of radius 0.005m with
velocity 0.3m/s.
1
Expert's answer
2019-08-05T12:49:54-0400

"Re=\frac{\rho\times v\times d}{\mu}"

Assume the liquid is water having density =1000kg/m3

Dynamic viscosity of water is 0.001003 kg/(m-s) at 20 degree C

velocity of flow is 0.3m/s

Diameter of pipe is 2x0.005=0.01m

Thus

Reynold no is (1000x0.30x0.01)/0.001003

Reynold no is found as 2,991.02692

Note if the Reynold no is less than 2000 then the flow is laminar

If the flow is more than 4000 then the flow is turbulent

If the flow is in transition then the Reynold no is in the range of 2000 to 4000

As the Reynold no found is 2,991.02692 so the flow is in transition range